For the reaction: C2H4Br2 + 3 KI → C2H4 + 2 KBr + KI3, when the rate of disappearance of KI is 6.0 × 10-5, what is the rate of appearance of C2H4?

a) 1.0 x 10^-5

b) 6.0 x 10^-5

c) 2.0 x 10^-5

d) 0.67 x 10^-5

e) 4.0 x 10^-5
Answer #1

rate = 1/3 d[KI]/dt = d[C2H4] /dt

1/3 d[KI]/dt = d[C2H4] /dt

d[KI]/dt = 6.0 x 10-5

d[C2H4] /dt = 1/3 (6.0 x 10-5)

d[C2H4] /dt = 2.0 x 10-5

answer = option c =  2.0 x 10-5